OpenStudy (arhin):

simplify 9a^2 - 4(a-b)^2

OpenStudy (anonymous):

it looks like a difference of squares x^2 - y^2 where x = 3a y = 2(a-b) good luck :)

OpenStudy (aonz):

(a+b)(a-b) = a^2 - b^2

OpenStudy (anonymous):

9a^2 - 4(a-b)^2 9a^2 - 4(a^2+b^2-2ab) 9a^2 - 4a^2-4b^2+8ab 5a^2-4b^2+8ab